The boss gives his employees a 5% pay rise.
Paul used to earn £8 per hour.
How much does he get after the rise?
Sandy used to earn a salary of £14300
How much did she get after the rise?

Answer by TimothyLamb(4379)

paul:
amount of rise: 8 * 0.05 = 0.40
hourly rate after rise: 8 * 1.05 = 8.40
---
sandy:
amount of rise: 14300 * 0.05 = 715
salary after rise: 14300 * 1.05 = 15015
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php